please help me with {x+6y=2
{5x+4y=36

Answer by elima(1433)

eqn.1 x+6y=2
eqn.2 5x+4y=36
We will begin this by eliminating one variable so we can solve the other.
Lets eliminate the x, to do that we need to multiply -5 to eqn.1
(-5)x+6y=2
-5x-30y=-10;
Now we will add eqn.1 and eqn.2 together;
-5x-30y=-10
5x+4y=36
-----------
-26y=26
y=-1
Now we have what y =, so we can plug that into either equation and find x;
x+6(-1)=2
x-6=2
x=8
Now we have our answer;
x=8, y=-1
Don't forget to check;
x+6y=2
8+6(-1)=2
8-6=2
Hope you understand
=)